OpenStudy (anonymous):

if x^2 > 0, then x < 0. Is that statement true or false?

OpenStudy (ivettef365):

what do you think ?

OpenStudy (anonymous):

x>0 or x<0

OpenStudy (anonymous):

x<0

OpenStudy (debbieg):

Can you find a counterexample?

OpenStudy (anonymous):

-1

OpenStudy (debbieg):

It says if x^2 > 0, then x < 0. So if you can find an x>0 such that if x^2 > 0, then that DISproves it, hence it is false.

OpenStudy (anonymous):

oh wait. nevermind -1*-1 = 1

OpenStudy (anonymous):

-1 fits the definition... what about something that works in the first part but not the second?

OpenStudy (debbieg):

Note, however, that the reverse is not true: finding an example for which the statement is true does not PROVE that it's true IN GENERAL. But finding an example that proves IT'S FALSE does prove that it's NOT TRUE in genera.

OpenStudy (anonymous):

2

OpenStudy (debbieg):

So what does that tell you? is it true or false?

OpenStudy (anonymous):

Pretty sure it's false. be> 0, or 4 >0, then 4 < 0???

OpenStudy (anonymous):

because 2^2 > 0

OpenStudy (debbieg):

Exactly. :)

OpenStudy (radar):

What if x has a complex value like 2i ?

OpenStudy (anonymous):

WOOOT THANKS DEBBIE

OpenStudy (radar):

x = -2, x^2 =4 x =2 , x^2 =4 x=2i, x^2=-4

OpenStudy (anonymous):

just no.

OpenStudy (anonymous):

but which is greater? 2i or -2i? also, is either of these greater or less than 0?

OpenStudy (anonymous):

greater

OpenStudy (anonymous):

can't compare complex numbers... only their moduli!

OpenStudy (anonymous):

so?..

OpenStudy (radar):

@pgpilot326 Equal but opposite. and when both are squared (-2i)^2 or (2i)^2 the result is still -4 which is less than zero. The drawing shows their relationship to zero, I will let you decide which one is less than or greater than zero :) l

OpenStudy (anonymous):

1+i or 1 - i?

OpenStudy (anonymous):

besides is 2i >0? if so, is -2i>0? can both be?

OpenStudy (anonymous):

not really?